### v1

# Adaptive Quantum Control

Abstract
This paper presents a model-free, learning-based adaptive controller for the density tracking problem in a two-level Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) quantum system. The adaptive controller is based on the continuous-time retrospective cost adaptive control. To preserve the geometric properties of the quantum system, an adaptive PID controller driven and optimized by Ulhmann’s fidelity is used. The proposed controller is validated in simulation for a low and a high-entropy density-tracking problem.
keywords: adaptive control, quantum control, density-tracking.
I INTRODUCTION
Quantum dynamical systems have several applications in sensing, metrology, communication, computation, and the internet [giovannetti2011advances, toth2014quantum, taylor2016quantum, cacciapuoti2020entanglement, nielsen2001quantum]. The reliable operation of quantum circuits, such as quantum logic gates, requires the precise and predictable behavior of quantum dynamical systems, much like how classical digital circuits depend on the accurate design of their fundamental electrical components, such as resistors and transistors [barenco1995elementary]. In quantum computing, the basic unit of information is the qubit, which, unlike classical bits limited to two distinct states (00 or $1$), can exist in a superposition of both states, enabling quantum algorithms to perform computations exponentially faster than classical systems [grover1996fast, shor1999polynomial]. The successful realization of quantum devices depends on the precise control of quantum dynamical systems.
Quantum control involves manipulating quantum systems to achieve specific outcomes. It has applications in fields such as quantum computing, communication, and sensing. This is achieved through external controls, such as electromagnetic fields, to direct the evolution of quantum systems. The nature of the desired quantum event in a control problem depends on the quantum problem objective. For instance, in applications such as quantum computing, metrology, communication, and optical systems [yamamoto2024engineering, zaletel2021preparation], a low entropy state is desired to reduce the randomness and uncertainties from the output, whereas high entropy is desired when high randomness is needed in quantum phenomena. Applications such as heat spin engines, quantum cryptography, and quantum thermodynamics simulation [von2019spin, Sehgal2021, somhorst2023quantum], can be conceived as quantum control problems of this category. Quantum control strategies often rely on optimizing the system’s time evolution, governed by quantum operators such as Hamiltonians and density matrices. The Schrödinger equation is used for closed systems, while open systems interacting with the environment follow the LGKS equation [brif2010control]. Recent advancements have improved control techniques for both isolated and open quantum systems [dong2010quantum, cong2014control]. However, due to fundamental physical limitations of quantum state measurement and uncertainty in the knowledge of quantum dynamics, the performance achievable by model-based control systems is limited.
Adaptive quantum control addresses the challenges posed by fundamental physical limitations in quantum state measurement and uncertainties in quantum dynamics. This control strategy adjusts its parameters based on feedback from the quantum system, compensating for variations in system parameters and external disturbances[Kosut2003, kosut2013adaptive, Liu2020]. By incorporating online learning algorithms, adaptive control reduces the impact of incomplete system knowledge and measurement limitations [egger2014adaptive] This approach is significant for practical quantum technologies, including quantum error correction and quantum information processing, where uncertainties affect performance.
Despite these advancements, existing adaptive controllers primarily address parameter uncertainty while still requiring some knowledge of the quantum system model, which can be difficult to obtain in complex quantum systems. This fundamental assumption enables adaptive controllers in the literature to focus mainly on managing parameter uncertainties [egger2014adaptive, Koswara2014, Hu2016].
This paper considers the problem of designing an adaptive controller for an open quantum system without any prior knowledge of the system’s dynamics. In this preliminary work, we consider Uhlmann’s fidelity as the error metrics [UHLMANN1976273]. Uhlmann’s fidelity is a unique real scalar metric that allows one to compare the similarity of two quantum states. In this paper, we design a continuous-time adaptive controller based on the Retrospective Cost Adaptive Control (RCAC) Technique [ali2015retrospective, Goel2015, Sobolic2016, poudel2023learning] for the tracking problem of an open quantum system whose dynamics are governed by the Lindblad-Gorini-Kossakowski-Sudarshan (LGKS) equation [manzano2020short]. The contributions of this paper are thus 1) the design of a continuous-time adaptive controller optimized by RCAC for the open quantum system described by the LGKS equation without prior knowledge of the system’s dynamics, 2) the design of an adaptive PID controller driven by the Uhlmann’s fidelity, and 3) numerical validation of the proposed controller in a low and high entropy state-tracking problem.
The paper is organized as follows. Section II describes the LGKS equation in the context of a controlled quantum open system. Section III presents the mathematical foundations for continuous time RCAC with forgetting factors. Section IV describes the numerical simulation that validate the application of the adaptive controller in a two-level open system. Finally, the paper concludes with a discussion of results and future research directions in section V.
II Problem Formulation
Consider the Lindblad-Gorini-Kossakowski-Sudarshan equation [manzano2020short]
$\dot{\rho}$ $=-i\hbar[H,\rho]+\sum_{i=1}^{n}\left(L_{i}\rho L_{i}^{\rm H}-\dfrac{1}{2}\{L_{i}^{\rm H}L_{i},\rho\}\right),$ (1)
where $\rho\in\mathbb{C}^{2\times 2}$ is the density matrix, the system Hamiltonian $H\in\mathbb{C}^{2\times 2}$ is Hermitian, $n$ is the number of damping terms, and for $i\in(1,2,\ldots,n),$ the matrices $L_{i}\in{\mathbb{C}}^{2\times 2}$ are the jump operators. Without loss of generality, the units are chosen such that $\hbar=1.$ The diagonal elements of $\rho$ represent the probability of a quantum event. Thus, in the case where $\rho\in{\mathbb{C}}^{2\times 2},$ (1) is called a two-level quantum dynamic system.
It is shown in [breuer2002theory, p 123] that, if ${\rm tr}\,\rho(0)=1,$ then, for all $t>0,$ $\rho(t)$ given by (1) satisfies ${\rm tr}\,(\rho(t))=1.$ Furthermore, for all $t\geq 0,$ $\rho(t)$ is positive semidefinite and Hermitian.
As shown in [altafini2012modeling], the Hamilotonian $H$ can be decomposed as
$H$ $=H_{0}+H_{1}u,$ (2)
where $H_{0}\in{\mathbb{C}}^{2\times 2}$ is the free Hamiltonian, $H_{1}\in{\mathbb{C}}^{2\times 2}$ is the control Hamiltonian, and $u$ is the control signal to be designed. In this paper, we consider the case where $u\in{\mathbb{R}},$ which ensures that the Hamiltonian $H$ is Hermitian at all time instants.
The control objective is to design an adaptive feedback control law such that the density matrix $\rho$ converges to the desired density matrix $\rho_{\rm d}.$
II-A Density Error
As shown in [jones2024controlling], the Uhlmann–Jozsa fidelity between two density matrices $\rho_{1},\rho_{2}$ is defined as
$F(\rho_{1},\rho_{2})$ $\stackrel{{\scriptstyle\triangle}}{{=}}\left({\rm tr}\,\left(\sqrt{\sqrt{\rho_{1}}\,\rho_{2}\,\sqrt{\rho_{1}}}\right)\right)^{2}.$ (3)
It follows from Theorem A.5 that $F(\rho_{1},\rho_{2})\in[0,1].$ Furthermore, $F(\rho_{1},\rho_{2})=1$ if and only if $\rho_{1}=\rho_{2}.$ Finally, as shown in [jozsa1994fidelity], in the case where $\rho_{1},\rho_{2}\in\mathbb{C}^{2\times 2},$ the Uhlmann–Jozsa fidelity is given by
$F(\rho_{1},\rho_{2})$ $={\rm tr}\,(\rho_{1}\rho_{2})+2\sqrt{\det(\rho_{1})\det(\rho_{2})}$ (4)
In this paper, we consider the density error
$e$ $\stackrel{{\scriptstyle\triangle}}{{=}}1-F(\rho,\rho_{\rm{d}})\in[0,1],$ (5)
where $\rho_{\rm{d}}\in\mathbb{C}^{2\times 2}$ is the desired density
II-B Adaptive PID Control
An adaptive proportional-integral-derivative controller can be written as
$u(t)$ $=k_{\rm p}(t)e(t)+k_{\rm i}(t)\int_{0}^{t}e(\tau){\rm d}\tau+k_{\rm d}(t)\dot{e}(t),$ (6)
where $k_{\rm p}(t),k_{\rm i}(t),$ and $k_{\rm d}(t)$ are the adaptive proportional, integral, and derivative gains. Note that the controller can be written as
$u(t)=\Phi(t)\theta(t),$ (7)
where
$\Phi(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}e(t)&\gamma(t)&\dot{e}(t)\end{bmatrix},\quad\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\begin{bmatrix}k_{\rm p}(t)\\ k_{\rm i}(t)\\ k_{\rm d}(t)\end{bmatrix}.$ (8)
Note that integral state $\gamma(t)$ satisfies $\dot{\gamma}(t)=e(t).$
The objective of the adaptive controller is to update the controller gains and thus compute the control signal $u(t)$ such that the density error $e(t)$ converges to zero. In this paper, we use the continuous-time retrospective cost adaptive control algorithm, described in the next section, to update the controller gains.
III Continuous Time Retrospective Cost Adaptive Control
Consider a dynamic system
$\dot{x}(t)$ $=f(x(t),u(t)),$ (9)
$y(t)$ $=g(x(t)),$ (10)
where $x(t)\in\mathbb{C}^{l_{x}}$ is the state, $u(t)\in\mathbb{C}^{l_{u}}$ is the input, $y(t)\in\mathbb{C}^{l_{y}}$ is the measured output, and the vector functions $f$ and $g$ are the dynamics and the output maps. Define the performance variable
$z(t)\stackrel{{\scriptstyle\triangle}}{{=}}y(t)-r(t),$ (11)
where $r(t)$ is the exogenous reference signal. The objective is to design an adaptive output feedback control law to ensure that $z(t)\to 0.$
Consider a linearly parameterized control law
$u(t)=\Phi(t)\theta(t),$ (12)
where the regressor matrix $\Phi(t)\in\mathbb{C}^{l_{u}\times l_{\theta}}$ contains the measured data and the vector $\theta(t)\in\mathbb{C}^{l_{\theta}}$ contains the controller gains to be optimized. Various linear parameterizations of MIMO controllers are described in [goel_2020_sparse_para].
Next, using (11), define the retrospective performance
$\hat{z}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}z(t)+\Phi_{\rm f}(t)\hat{\theta}(t)-u_{\rm f}(t),$ (13)
where $\hat{\theta}(t)$ is the controller gain to be optimized and the filtered regressor $\Phi_{\rm f}(t)$ and the filtered control $u_{\rm f}(t)$ are defined as
$\Phi_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[\Phi(t)\right]\in\mathbb{C}^{l_{y}\times l_{\theta}},$ (14)
$u_{\rm f}(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}G_{\rm f}(s)\left[u(t)\right]\in\mathbb{C}^{l_{y}\times l_{\theta}},$ (15)
where $G_{\rm f}(s)$ is a dynamic filter.
The filtered signals $\Phi_{\rm f}(t)$ and $u_{\rm f}(t)$ are computed in the time domain, as shown below. Let $(A_{\rm f},B_{\rm f},C_{\rm f},D_{\rm f})$ be a realization of $G_{\rm f}(s).$ Then,
$\dot{x}_{\Phi}$ $=A_{\rm f}x_{\phi}+B_{\rm f}\Phi,$ (16)
$\Phi_{\rm f}$ $=C_{\rm f}x_{\phi}+D_{\rm f}\Phi,$ (17)
and
$\dot{x}_{u}$ $=A_{\rm f}x_{u}+B_{\rm f}u,$ (18)
$u_{\rm f}$ $=C_{\rm f}x_{u}+D_{\rm f}u.$ (19)
Next, define the retrospective cost
$J(t,\hat{\theta})$ $=\int_{0}^{t}\Big{(}e^{-\lambda(t-\tau)}\hat{z}(\tau)^{\rm T}R_{z}\hat{z}(\tau)$
$\quad+(\Phi(\tau)\hat{\theta})^{\rm T}R_{u}(\Phi(\tau)\hat{\theta})\Big{)}{\rm d}\tau$
$\quad+e^{-\lambda t}\hat{\theta}^{T}R_{\theta}\hat{\theta},$ (20)
where $R_{z},R_{u},$ and $R_{\theta}$ are positive definite weighting matrices of appropriate dimensions, and $\lambda<0$ is the exponential forgetting factor.
Proposition III.1.
Consider the cost function $J(t,\hat{\theta})$ given by (20). For all $t\geq 0,$ define the minimizer of $J(t,\hat{\theta})$ by
$\theta(t)\stackrel{{\scriptstyle\triangle}}{{=}}\underset{\hat{\theta}\in{\mathbb{R}}^{l_{\theta}}}{\operatorname{argmin}}J(t,\hat{\theta}).$ (21)
Then, for all $t\geq 0,$ the minimizer satisfies
$\dot{\theta}$ $=-P\Phi_{\rm f}^{\rm T}R_{z}(z+\Phi_{\rm f}\theta-u_{\rm f})-P\Phi^{\rm T}R_{u}\Phi\theta,$ (22)
$\dot{P}$ $=\lambda P-P\left(\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi\right)P$ (23)
where $P(0)=R_{\theta}^{-1}$ and $\theta(0)=0.$
Proof.
Note that the cost function (20) can be written as
$J(t,\hat{\theta})$ $=\hat{\theta}^{T}A(t)\hat{\theta}+2\hat{\theta}^{T}b(t)+c(t),$ (24)
where
$A(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\int_{0}^{t}e^{-\lambda(t-\tau)}\left(\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi\right){\rm d}\tau+e^{-\lambda t}R_{\theta},$
$b(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\int_{0}^{t}e^{-\lambda(t-\tau)}\Phi_{\rm f}^{\rm T}R_{z}(z-u_{\rm f}){\rm d}\tau,$
$c(t)$ $\stackrel{{\scriptstyle\triangle}}{{=}}\int_{0}^{t}e^{-\lambda(t-\tau)}(z-u_{\rm f})^{\rm T}R_{z}(z-u_{\rm f}){\rm d}\tau.$
The matrices $A(t)$ and $b(t)$ satisfy
$\dot{A}$ $=-\lambda A+\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi,$
$\dot{b}$ $=-\lambda b+\Phi_{\rm f}^{\rm T}R_{z}(z-u_{\rm f}),$
where $A(0)=R_{\theta}$ and $b(0)=0.$
Next, define, for all $\geq 0,$ $P(t)\stackrel{{\scriptstyle\triangle}}{{=}}A(t)^{-1}.$ Using the fact that $P(t)A(t)=I,$ it follows that
$\dot{P}(t)=-P(t)\dot{A}(t)P(t),$
and thus
$\dot{P}=\lambda P-P\left(\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi\right)P.$
Finally, note that the minimizer of (24) is given by
$\theta(t)=-A(t)^{-1}b(t)=-P(t)b(t),$
and thus
$\dot{\theta}(t)$ $=-\dot{P}b-P\dot{b}$
$=-P\left(\Phi_{\rm f}^{\rm T}R_{z}\Phi_{\rm f}+\Phi^{\rm T}R_{u}\Phi\right)PP^{-1}\theta$
$\quad-P\Phi_{\rm f}^{\rm T}R_{z}(z-u_{\rm f})$
$=-P\Phi_{\rm f}^{\rm T}R_{z}(z+\Phi_{\rm f}\theta-u_{\rm f})-P\Phi^{\rm T}R_{u}\Phi\theta.$
The control law is thus
$u(t)=\Phi(t)\theta(t),$ (25)
where $\theta(t)$ is given by (22).
IV Numerical Simulation
In this paper, we consider a two-level system, with dynamics given by the LGKS (1), with a single damped term, that is, $n=1$. The control system architecture is shown in Figure 1.
$F(\rho,\rho_{\rm{d}})$$\begin{array}[]{c}\text{Adaptive}\\ \text{Controller}\end{array}$ $\mathcal{L}\mathcal{G}\mathcal{K}\mathcal{S}$$\rho_{\rm{d}}$$e$$u$$\rho$
Figure 1: Adaptive control architecture to track desired density matrices in the LGKS system.
To simulate the LKGS equation, we set
$H_{0}=0.5\begin{bmatrix}1&0\\ 0&-1\end{bmatrix},H_{1}=0.5\begin{bmatrix}0&1\\ 1&0\end{bmatrix},L_{1}=\begin{bmatrix}0&1\\ 0&0\end{bmatrix},$ (26)
where $H_{0}$ and $H_{1}$ are Pauli matrices that, as described in [d2001optimal, kamakari2022digital], satisfy the controllability conditions in the problem of controlling a two-level quantum system, and $L_{1}$ is selected as a non-Hermitian matrix, as suggested in [verstraete2009quantum, schlimgen2022quantum].
The initial density matrix $\rho(0)$ is assumed to be
$\rho_{0}$ $=\begin{bmatrix}0.4&0.1+0.3\imath\\ 0.1-0.3\imath&0.6\end{bmatrix}.$ (27)
Note that $\rho(0)$ is randomly generated such that ${\rm tr}\,\rho(0)=1$ and it is Hermitian. As shown in [moses2015creation] and [xu2021multicriticality], we consider the problem of moving the quantum state $\rho(0)$ to a low and a high desired entropy states $\rho_{\rm d},$ where the Von Neumann entropy of a quantum state $\rho$ is given by
$S(\rho)\stackrel{{\scriptstyle\triangle}}{{=}}-{\rm tr}\,(\rho\,\rm{ln}(\rho)).$ (28)
First, we consider the problem of driving the quantum state $\rho(t)$ to
$\rho_{\rm d}=\begin{bmatrix}0.8571&0.2857+0.1429\imath\\ 0.2857-0.1429\imath&0.1429\end{bmatrix}.$ (29)
Note that $S(\rho_{\rm d})=0.1013,$ and thus, this desired state is a low-entropy state. Furthermore, note that $\rho_{\rm d},$ given by (29) is an equilibrium point of (1) in the case where $u(t)\equiv 1.$
With the desired density $\rho_{\rm d}$ given by (29) as the reference for the controller, we first tune the RCAC hyperparameters. In particular, in RCAC, we set $R_{z}=1,$ $R_{u}=1,$ and $\lambda=0.01$; and optimize $P_{0}$ and the filter coefficient $\beta,$ where $G_{\rm f}(s)=\dfrac{1}{s+\beta},$ as described below. We optimize only $P_{0}$ and the filter coefficient since it has been observed in several numerical experiments that the choice of $P_{0}$ and the filter coefficient has the most dominant impact on the closed-loop performance. For each pair of $(P_{0},\beta)\in\{(10^{-5},10^{-4},\ldots,10^{10})\times(0,1,2,5,100,2000)\},$ we apply the RCAC algorithm to update the adaptive PID controller to drive the quantum state from $\rho_{0}$, given by (27) to $\rho_{\rm d}$, given by (29). Each simulation is run for 200 seconds. At the end of each simulation, the hyperparameter cost
$J_{h}\stackrel{{\scriptstyle\triangle}}{{=}}\int_{190}^{200}|z(\tau)|{\rm d}\tau$ (30)
is computed. Figure 2 shows the contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The tuning pair with the minimum value of $J_{h}$ is selected for all the numerical simulations in the rest of the paper. In particular, in all the examples, we set $P_{0}=10^{-3}\times I_{3},\beta=2000.$
Figure 2: Contour plot of $J_{h}$ for various values of $(P_{0},\beta).$ The pair $(P0,\beta)$ that minimizes $J_{h}$ is used in all numerical simulations.
IV-A Low-entropy Density
With the selected RCAC hyperparameters, Figure 3 shows the absolute value of the density error $e(t)$ between the system state and the desired low-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ given by adaptive controller. Figure 4 shows the real and imaginary part of the quantum state $\rho(t)$ and the Von Neumann entropy of the quantum state. Figure 5 shows the trajectory of the quantum state of the quantum system (1) on the Bloch sphere. Note that the quantum state $\rho$ satisfies $x=2\,\rm{Real}(\rho_{1,2}),$ $y=2\,\rm{Imag}(\rho_{1,2}),$ and $z=\rho_{1,1}-\rho_{2,2}$ on the Bloch sphere [wie2020two].
Figure 3: Low-entropy density tracking. The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ given by adaptive controller.
Figure 4: Low-entropy density tracking. The first two subplots show the real and imaginary parts of the quantum state $\rho(t).$ Note that the desired state component is shown in dashes and the system response is shown in solid with the corresponding color. The third subplot shows the entropy of the quantum state.
Figure 5: Low-entropy density tracking. Trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
IV-B High-entropy Density
Next, we consider the problem of driving the quantum state $\rho(t)$ to
$\rho_{\rm d}=\begin{bmatrix}0.5168&0.0971+0.0460i\\ 0.0971-0.0460i&0.4832\end{bmatrix}.$ (31)
Note that $S(\rho_{\rm d})=0.6693,$ and thus, this desired state is a high-entropy state. Furthermore, note that $\rho_{\rm d},$ given by (31) is an equilibrium point of (1) in the case where $u(t)\equiv 10.$
Figure 6 shows the absolute value of the density error $e(t)$ between the system state and the desired high-entropy state on a logarithmic scale, the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the control $u$ given by adaptive controller. Figure 7 shows the real and imaginary part of the quantum state $\rho(t)$ and the Von Neumann entropy of the quantum state. Figure 8 shows the trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
The two examples described above suggest that the adaptive PID controller updated by the RCAC algorithm can effectively track an arbitrary desired density matrix $\rho_{\rm d}$ without requiring any model of the underlying dynamics of the density matrix.
Figure 6: High-entropy density tracking. The first subplot shows the absolute value of the density error $e(t)$ between the system state and the desired state on a logarithmic scale, the next three subplots show the adaptive gains $k_{\rm p},$ $k_{\rm i},$ and $k_{\rm d},$ updated by RCAC, and the last subplot shows the control $u$ given by adaptive controller.
Figure 7: High-entropy density tracking. The first two subplot shows the real and imaginary part of the quantum state $\rho(t).$ Note that the desired state component is shown in dashes and the system response is shown in solid with the corresponding color. The third subplot shows the entropy of the quantum state.
Figure 8: High-entropy density tracking. Trajectory of the quantum state of the quantum system (1) on the Bloch sphere.
V Conclusions
This paper presented an online learning-based, model-free adaptive controller for tracking a desired density matrix in a two-level Lindblad-Gorini-Kossakowski-Sudarshan quantum system. In particular, the adaptive controller uses the retrospective cost adaptive control algorithm. Since the classical RCAC was developed for a discrete-time system, the paper presented a continuous-time extension of RCAC. An adaptive PID controller structure was chosen to generate a control signal that preserves the LKGS solution’s geometric structure. Uhlmann-Jozsa fidelity was used as an error metric to drive the adaptive PID controller. Numerical simulations show that the proposed controller effectively tracks desired low and high-entropy states without requiring a model of the underlying system dynamics.
Future work will extend the technique to the case where the input is a Hermitian matrix and the density coherence is used as an error metric. This will allow higher-dimensional parameterization of the adaptive PID controller, which may yield better tracking performance.
[…]
Appendix
A Some Useful Facts
This appendix reviews some useful definitions and facts about Hermitian matrices used in the LKGS system.
Definition A.1.
Let $A\in{\mathbb{C}}^{n\times n}.$ $A$ is a Hermitian matrix if $A=\overline{A^{\rm T}},$ where $\overline{A^{\rm T}}$ is the conjugate transpose of $A.$
The conjguate transpose of $A$ is denoted by $A^{\rm H}.$ If $A$ is Hermitian, then $A=A^{\rm H}.$
Definition A.2.
Let $A,B\in{\mathbb{C}}^{n\times n}.$ Then, the commutator $[A,B]$ is defined as
$[A,B]\stackrel{{\scriptstyle\triangle}}{{=}}AB-BA,$ (32)
and the anticommutator $\{A,B\}$ is defined as
$\{A,B\}\stackrel{{\scriptstyle\triangle}}{{=}}AB+BA.$ (33)
Fact A.1.
Let $A,B\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $-\imath[A,B]$ is Hermitian.
Proof.
Note that $(-i[A,B])^{\rm H}=(-\imath)^{\rm H}[A,B]^{\rm H}=\imath\left(AB-BA\right)^{\rm H}=\imath(A^{\rm H}B^{\rm H}-B^{\rm H}A^{\rm H})=\imath(BA-AB)=-\imath(AB-BA)=-\imath[A,B]$. Thus, $-\imath[A,B]$ is Hermitian. ∎
Fact A.2.
Let $U\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $U^{\rm H}U$ and $UU^{\rm H}$ are Hermitian.
Proof.
Note that $(U^{\rm H}U)^{\rm H}=U^{\rm H}(U^{\rm H})^{\rm H}=U^{\rm H}U,$ which implies that $UU^{\rm H}$ is Hermitian. A similar argument proves that $U^{\rm H}U$ is also Hermitian. ∎
Fact A.3.
Let $A,U\in{\mathbb{C}}^{n\times n}$ be Hermitian. Then, $UAU^{\rm H}$ is Hermitian.
Proof.
Note that $(UAU^{\rm H})^{\rm H}=(U^{\rm H})^{\rm H}(UA)^{\rm H}=UA^{\rm H}U^{\rm H}=UAU^{\rm H},$ which implies that $UAU^{\rm H}$ is Hermitian. ∎
Fact A.4.
Let $A,B,C\in{\mathbb{C}}^{n\times n}$ be Hermitian. Suppose $B$ satisfies
$\dot{B}=-\imath[A,B]+\sum_{i=1}^{N}\left(CBC^{\rm H}-\dfrac{1}{2}\{C^{\rm H}C,B\}\right).$ (34)
Then, for all $t\geq 0,$ ${\rm tr}\,B$ is a constant.
Proof.
Note that $\dfrac{{\rm d}}{{\rm d}t}{\rm tr}\,(B)={\rm tr}\,(\dot{B})$. Next, note that ${\rm tr}\,(-\imath[A,B])=-\imath{\rm tr}\,(AB-BA)=-\imath{\rm tr}\,(AB)+\imath{\rm tr}\,(BA)=-\imath{\rm tr}\,(AB)+\imath{\rm tr}\,(AB)=0.$ Using the cyclic property of the trace, it follows that
${\rm tr}\,\left(\sum_{i}^{N}\left(CBC^{\rm H}-\dfrac{1}{2}(C^{\rm H}CB+BC^{\rm H}C)\right)\right)$
$=\sum_{i}^{N}{\rm tr}\,(CBC^{\rm H})-\dfrac{1}{2}\sum_{i}^{N}{\rm tr}\,(C^{\rm H}CB)-\dfrac{1}{2}\sum_{i}^{N}{\rm tr}\,(C^{\rm H}CB)$
$=\sum_{i}^{N}{\rm tr}\,(C^{\rm H}CB)-\sum_{i}^{N}{\rm tr}\,(C^{\rm H}CB)=0,$
which completes the proof. ∎
Fact A.5.
Let $\rho\in\mathbb{C}^{n\times n}$ and $\sigma\in\mathbb{C}^{n\times n}$ be two Hermitian PSD matrices with traces equal to $1$. Then,
$1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\sigma\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0$
Proof.
Since $\rho$ and $\sigma$ are PSD Hermitian. Then, $\rho^{{\rm H}}=\rho$, $\sigma^{{\rm H}}=\sigma$, $x^{{\rm H}}\rho x\geq 0$ and $x^{{\rm H}}\sigma x\geq 0$ for all $x\in\mathbb{C}.$ Furthermore, the unique square root of a PSD Hermitian matrix is also Hermitian and PSD [horn2012matrix, p 440]. This implies that, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\left(\sqrt{(}\rho)\sigma\right)^{{\rm H}}=\sqrt{\rho}^{{\rm H}}\sigma^{{\rm H}}\sqrt{\rho}^{{\rm H}}=\sqrt{\rho}\sigma\sqrt{\rho}.$ Define, $y\stackrel{{\scriptstyle\triangle}}{{=}}\sqrt{\rho}x\in\mathbb{C}^{n}.$ Then, $x^{{\rm H}}\sqrt{\rho}\sigma\sqrt{\rho}x=y^{{\rm H}}\sigma y\geq 0,$ which implies that $\sqrt{\rho}\sigma\sqrt{\rho}\geq 0.$ Therefore, $\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)$ is Hermitian, and $\sqrt{\left(\sqrt{(}\rho)\sigma\sqrt{\rho}\right)}$ is Hermitian and unique. From the Cauchy-Schwarz inequality and the cyclic property for traces, $|{\rm tr}\,(\sqrt{\sqrt{\rho}\sigma\sqrt{\rho}})|^{2}=|{\rm tr}\,(\sqrt{\sqrt{\rho}\sqrt{\rho}\sigma})|^{2}=|{\rm tr}\,(\sqrt{\rho\sigma})|^{2}\leq{\rm tr}\,(\rho){\rm tr}\,(\sigma)\leq 1$. Thus, $1\geq\left({\rm tr}\,\left(\sqrt{\sqrt{\rho}\,\rho_{\rm{d}}\,\sqrt{\rho}}\right)\right)^{2}\in\mathbb{R}\geq 0$